OpenStudy (anonymous):

Calculate how many grams of oxygen form when each quantity of reactant completely reacts. 2HgO----->2Hg+O2

OpenStudy (anonymous):

since you dont have amount of any reactant/product lets take you have 100 g of HgO and go from there: n(O2)/n(HgO) = 1/2 --> n(O2)= 1/2 * n(HgO) n(O2) = 1/2 * m(HgO)/M(HgO) m(O2)= 1/2 * m(HgO)/M(HgO) * M(O2)